OpenStudy (anonymous):

Solve this problem. A man is 65 years old and his grandson is 20. How many years ago was the grandfather six times as old as the grandson?

OpenStudy (anonymous):

(65 - x) / (20 / x) = 6 65 - x = 120 - 6x 5x = 55 x = 11

OpenStudy (anonymous):

65-20 = 45 45/5=9 *i used 5 because it 6 times but the first time is the grandsons age... so 11 years ago